(x^2D^2-xD+1)y=sin (logx)

1
Expert's answer
2021-12-14T13:39:32-0500

"Given\mathrm{:}\ \ \ \left(x^{\mathrm{2}}D^{\mathrm{2}}\ \ \ -\ \ xD\ \ +\ \ \mathrm{1}\right)y\ \ =\ \ \mathrm{sin}\left(\mathrm{log}\left(x\right)\right) \\ 
Let\ \ z=\mathrm{log}\left(x\right)\ \ ,\ \ such\ \ that\ \ x=e^z \\ 
 \\ 
D\left(D-\mathrm{1}\right)y-Dy\ \ +\ \ y\ \ =\ \ \mathrm{sin}\left(z\right) \\ 
 \\ 
\left(D^{\mathrm{2}}-\mathrm{2}D+\mathrm{1}\right)y\ \ =\ \ \mathrm{sin}\left(z\right) \\ 
 \\ 
A.E.\ \ is\ \ \ \ m^{\mathrm{2}}-\mathrm{2}m{}{}\ +\ \mathrm{1}=0\ \ \ \\ 
 \\ 
\Rightarrow \ \ \left(m-\mathrm{1}\right)\left(m-\mathrm{1}\right)=0\ \ \ ,\ \ m=\mathrm{1,1}\ \ \ \ \ \left(repeated\ \ root\right) \\ 
 \\ 
\Rightarrow y_c=\left(A+Bz\right)e^z\ \ ...................\left(\mathrm{1}\right) \\ 
 \\ 
Let\ \ y_p=\frac{\mathrm{1}}{D^{\mathrm{2}}-\mathrm{2}D+\mathrm{1}}\mathrm{sin}\left(z\right) \\ 
 \\ 
y_p=\frac{\mathrm{1}}{-{\mathrm{1}}^{\mathrm{2}}-\mathrm{2}D+\mathrm{1}}\mathrm{sin}\left(z\right)\ \ =\frac{\mathrm{1}}{-\mathrm{2}D}\mathrm{sin}\left(z\right)\ \ =\frac{\mathrm{cos}\left(z\right)}{\mathrm{2}} \\ 
 \\ 
y_p=\frac{\mathrm{cos}\left(z\right)}{\mathrm{2}}..........................\left(\mathrm{2}\right) \\ 
 \\ 
y=y_c+y_p \\ 
 \\ 
y=\left(A+Bz\right)e^z\ +\frac{\mathrm{cos}\left(z\right)}{\mathrm{2}} \\ 
 \\ 
y=\left(A+B\mathrm{log}\left(x\right)\right)x\ +\frac{\mathrm{cos}\left(\mathrm{log}\left(x\right)\right)}{\mathrm{2}}"